OpenStudy (anonymous):

write an equation in slope intercept form for the line that satisfies the following conditions x intercept 12 and y intercept 14

OpenStudy (karatechopper):

if they are telling you the x intercept and y intercept you have your two points which will lead you into the problem! x intercept: 12 y intercept: 14 x intercept point: (12,0) y intercept point: (0,14) Now does this make the question easier?:) Welcome to OpenStudy, btw! :P

OpenStudy (zzr0ck3r):

y=mx+b is the form you want m = slope b = y intercept so you know it looks like this y = mx+14 we must find what m is, so we use the point they gave you for the x intercept (12,0) so make x = 12 and y = 0 0 = m(12)+14 can you solve for m?

OpenStudy (anonymous):

equation of the line with x-intercept 'a' and y-intercept 'b' is given by x/a + y/b =1 here a=12 and b=14

OpenStudy (anonymous):

so check x/12 +y/14 =1 7x+6y=84

OpenStudy (anonymous):

The eq of line in double intercept form is given as below \[\frac{x}{a} + \frac{y}{b}=1\] where a= x intercept = 12 and b= y intercept = 14 \[\frac{x}{12} + \frac{y}{14}=1\] i.e. 14x+ 12y= 168 which is the required eq of the line.